If the speed of a transverse wave on a stretched string of length 1 m is 60 m/s what is the
fundamental frequency of vibration?

Grade:12

2 Answers

Arun
6 years ago
I = 1m, V = 60 m/s ∴ fundamental frequency, f base o = v/2; = 30 sec^1 = 30 Hz (answer)
Hope it helps.
vineet jakhar
5 years ago
v = 60 m/s l=1mfrequency(f) = {nv/2l} " where n is integer" therefore fundamental frequency will be - {1*60/2*1}=30 hz